### v2

# Can Agents Fix Agent Issues?

Abstract
LLM-based agent systems are emerging as a new software paradigm and have been widely adopted across diverse domains such as medicine, robotics, and programming. However, maintaining these systems requires substantial effort, as they are inevitably prone to bugs and continually evolve to meet changing external requirements. Therefore, automatically resolving agent issues (i.e., bug reports or feature requests) is a crucial and challenging task. While recent software engineering (SE) agents (e.g., SWE-agent) have shown promise in addressing issues in traditional software systems, it remains unclear how effectively they can resolve real-world issues in agent systems, which differ significantly from traditional software. To fill this gap, we first manually analyze 201 real-world agent issues and identify common categories of agent issues. We then spend 500 person-hours constructing AgentIssue-Bench, a reproducible benchmark comprising 50 agent issue resolution tasks (each with an executable environment and failure-triggering tests). We further evaluate state-of-the-art SE agents on AgentIssue-Bench and reveal their limited effectiveness (i.e., with only 3.33% - 12.67% resolution rates). These results underscore the unique challenges of maintaining agent systems compared to traditional software, highlighting the need for further research to develop advanced SE agents for resolving agent issues. Data and code are available at https://alfin06.github.io/AgentIssue-Bench-Leaderboard/#/.
1 Introduction
LLM-based agent systems have seen widespread adoption across diverse domains, such as medicine DBLP:journals/corr/abs-2502-18864 , programming devin ; DBLP:journals/corr/abs-2405-15793 , robotics DBLP:conf/iros/KannanVM24 ; DBLP:journals/corr/abs-2311-07226 , psychology qiu2024interactive , and general-purpose personal assistants crewai ; agixt . Driven by rapid advancements, agent systems are emerging as a new software paradigm, playing an increasingly pervasive role in shaping and supporting the full spectrum of human activities.
As products of human intellectual labor, similar as traditional software systems, agent systems are also inevitably prone to quality issues. Recent studies DBLP:journals/corr/abs-2503-13657 have shown that multi-agent systems exhibit diverse failure modes during operation. Moreover, agent systems are continuously evolving to meet changing external requirements, making their maintenance both crucial and labor-intensive. For instance, by May 2025, the agent system MetaGPT metagpt had accumulated over 800 GitHub issues (an issue is typically a bug report or a feature request), highlighting the substantial maintenance workload associated with agent systems.
Automating the issue resolution process has been an important and challenging direction with substantial dedicated research effort. In particular, with the recent advances in agent systems, there is a growing trend toward developing software engineering agents DBLP:journals/corr/abs-2407-01489 ; DBLP:conf/issta/0002RFR24 ; DBLP:journals/corr/abs-2405-15793 ; devin ; opendevin ; aider ; moatless (referred to as SE agents in this paper), which can automatically resolve real-world software issues. Recent SE agents have demonstrated strong potential in resolving issues in traditional software systems. For instance, Agentless DBLP:journals/corr/abs-2407-01489 correctly resolves 50.80% of issues on SWE-bench sweboard , a real-world issue resolution benchmark for traditional Python software.
Although SE agents have shown promise in resolving issues in traditional software systems, it remains unclear how effectively they perform on agent systems, which is a new software paradigm that differs significantly from traditional software. Therefore, in this work, we aim to answer the central question: can SE agents fix issues in agent systems?
To understand issues in agent systems, we first perform an empirical study to analyze and catalog real-world agent issues. In particular, we collect 201 real-world GitHub issues along with developer-committed patches from 16 widely-used agent systems. We further build a taxonomy of agent issues with human annotators via grounded theory, resulting in 6 categories and 20 sub-categories of common agent issues. Our taxonomy reveals that real-world agent systems exhibit a diverse range of issues, many of which possess unique characteristics not typically found in traditional software systems. The findings highlight the large engineering effort for maintaining agent systems, confirming that automated issue resolution for agent systems is a challenging and critical problem.
We then build AgentIssue-Bench, the first reproducible benchmark for agent issue resolution. Reproducing agent issues is particularly more challenging compared to traditional software issues, largely due to the nondeterminism of LLMs and the volatility of external resources (e.g., tools) that agents interact with. As a result, from the 201 issues analyzed, we invested 500 person-hours to successfully reproduce 50 agent issues. Each issue resolution task in AgentIssue-Bench is packaged within an executable Docker environment, along with failure-triggering tests, user-reported issue descriptions, the buggy version, and the developer-committed patched version of the codebase.
We further evaluate multiple state-of-the-art SE agents (i.e., Agentless DBLP:journals/corr/abs-2407-01489 , AutoCodeRover DBLP:conf/issta/0002RFR24 , and SWE-agent DBLP:journals/corr/abs-2405-15793 ) with both GPT-4o GPT4o and Claude-3.5-Sonnet claude on AgentIssue-Bench. We find that all of the existing SE agents exhibit limited capabilities in resolving agent issues. For instance, only 3.33% to 12.67% of agent issues are correctly resolved, which is significantly lower than the resolution rates achieved when these SE agents are applied to traditional software (e.g., 23.20% - 50.80% resolution rate sweboard ). We further conduct a qualitative analysis to break down the resolution capabilities of SE agents across different categories. Notably, the majority of resolved issues pertain to utility or dependency issues, while the most of LLM-related issues (e.g., compatibility with LLM providers or LLM operation issues) remain unsolved. Overall, our analysis reveals the limitations of current SE agents in resolving agent issues, underscoring the need for building advanced SE agents tailored to the maintenance of agent systems.
In summary, this work makes the following contributions:
  • Taxonomy. We present the first taxonomy of issues in agent systems, derived from extensive manual analysis, which summarizes the common maintenance demands encountered during agent system evolution.
  • Reproducible benchmark AgentIssue-Bench. We manually construct the first issue resolution benchmark of real-world agent issues. Each task is packed into an executable Docker environment, including issue descriptions, failure-triggering tests, and both buggy and patched versions of the codebase, enabling easy reproduction and validation through one-click execution.
  • Evaluation. We evaluate state-of-the-art SE agents on AgentIssue-Bench with both quantitative and qualitative analysis, and find their limited capabilities in solving agent issues. Our findings highlight the unique challenges of maintaining agent systems, underscoring the need to develop more powerful SE agents for resolving agent issues.
2 Background and Related Work
2.1 LLM-based Agent Systems
LLM-based agent systems are emerging as a new software paradigm, which have been widely applied across various fields (e.g., medicine DBLP:journals/corr/abs-2502-18864 , programming devin ; DBLP:journals/corr/abs-2405-15793 , robotics DBLP:conf/iros/KannanVM24 ; DBLP:journals/corr/abs-2311-07226 , psychology qiu2024interactive , and general-purpose personal assistants crewai ; agixt ) with remarkable abilities. An LLM-based agent system DBLP:journals/chinaf/XiCGHDHZWJZZFWXZWJZLYDW25 ; DBLP:journals/fcsc/WangMFZYZCTCLZWW24 typically consists of: (i) an LLM-controlled brain that decomposes and schedules tasks (i.e., planning) and records the historical behaviors (i.e., memory); (ii) a perception component that receives information from the environment; and (iii) an action component that interacts with the environment by invoking external tools. In addition, single-agent systems can collaborate to form multi-agent systems, which can tackle more complex tasks with better flexibility and effectiveness.
Quality problems in LLM-integrated systems. Given the widespread adoption of LLMs, recent work has been looking into quality problems (e.g., bugs or runtime failures) in LLM-integrated systems. For example, Shao et al. shao2025llms catalog the integration bugs in LLM-integrated systems. Different from their work, our work focuses on a specific category of LLM-integrated systems, i.e., LLM-based agent systems. Along this direction, Cemri et al. DBLP:journals/corr/abs-2503-13657 build a taxonomy of failure modes in multi-agent systems. While their work focuses on runtime failure symptoms by analyzing failure trajectories, our taxonomy centers on agent issue resolution by analyzing both real-world user-reported issues and developer-committed patches. Therefore, our work complements existing efforts by providing a perspective on maintaining agent systems, encompassing a broader scope that includes not only bug fixes but also feature requests. Moreover, our work is further different from existing work by introducing the first reproducible benchmark for agent issue resolution and empirically evaluating state-of-the-art SE agents on their ability to resolve agent issues.
2.2 Software Engineering Agents
Software Engineering (SE) agents are a category of agent systems specifically designed to tackle SE tasks DBLP:journals/corr/abs-2409-02977 . In particular, there is a growing trend in both industry and academia toward developing SE agents devin ; opendevin ; DBLP:journals/corr/abs-2407-01489 ; DBLP:conf/issta/0002RFR24 ; moatless ; DBLP:journals/corr/abs-2405-15793 ; aider ; DBLP:journals/corr/abs-2403-16362 ; DBLP:journals/corr/abs-2409-19894 , which can support end-to-end software maintenance by automatically resolving user-reported issues (e.g., bug fixes or feature requests). For instance, Devin devin is one of the first SE agents capable of resolving software issues by invoking file editors, terminals, and search tools. More recently, SWE-agent DBLP:journals/corr/abs-2405-15793 interacts with the code repository environment through a custom Agent-Computer Interface (ACI), capable of performing actions such as manipulating files and executing bash commands; AutoCodeRover DBLP:conf/issta/0002RFR24 incorporates a suite of code search tools that iteratively retrieve relevant code contexts to navigate the repository and localize issue locations; Moatless moatless equips agents with code search and retrieval tools to identify the issue locations; Agentless DBLP:journals/corr/abs-2407-01489 optimizes the agent workflow with human expertise, incorporating hierarchical localization and regression testing to improve issue resolution rates. In this work, we evaluate the effectiveness of existing SE agents in resolving issues in agent systems.
Benchmarking issue resolution capabilities of SE agents. With the rise of SE agents, an increasing number of benchmarks have been developed to evaluate their capabilities in addressing real-world issue resolution tasks. For instance, Jimenez et al. swebench build SWE-bench from GitHub issues of 12 Python libraries. Based on SWE-bench, researchers further propose a series of benchmarks, e.g., SWE-bench Lite swebench , SWE-bench verified sweverified , and SWE-bench Lite-S DBLP:journals/corr/abs-2407-01489 , which are refined versions of SWE-bench with additional quality checking. While the SWE-bench series only includes issues of Python software, Zan et al. DBLP:journals/corr/abs-2408-14354 ; DBLP:journals/corr/abs-2502-12115 further propose SWE-bench Java, an issue resolution benchmark for Java software, and Yang et al. DBLP:conf/iclr/YangJZLYWPMSNY025 build SWE-bench Multimodal, comprising frontend issue resolutions tasks from open-source JavaScript libraries. More recently, OpenAI releases SWE-Lancer Diamond swelancer , an issue resolution benchmark with end-to-end tests for both open-source and commercial Expensify expensify software. While existing benchmarks focus exclusively on issue resolution in traditional software systems, our work introduces the first reproducible benchmark targeting issues in agent systems, an emerging software paradigm with features distinct from traditional software. Using this benchmark, we find that current SE agents are still unable to resolve the majority of issue resolution tasks in agent systems.
3 Agent Issue Taxonomy
To understand issues during agent system maintenance, we first manually analyze and categorize real-world GitHub issues in widely-used agent systems.
3.1 Methodology
Figure 1 illustrates our methodology of systematically collecting and analyzing agent issues.
3.1.1 Data Collection
Agent system collection. To select diverse and representative agent systems, we first use the GitHub search API to obtain 50 repositories with keywords “AI agents” by Feb 2025. We then manually go through each repository to keep the ones that are LLM-based agent systems (filter out the unrelated ones like paper lists or tutorials); to focus on agent systems with active maintenance, we only keep the ones with more than 1k stars and 30 issues. In this way, we collect 16 agent systems, such as MetaGPT metagpt , AutoGen autogen , GPT-engineer gpt-engineer , and CrewAI crewai . The full list of our analyzed agent systems is in Appendix A.
Figure 1: Methodology of agent issue taxonomy construction.
Agent issue extraction. For each studied agent system, we adopt the following inclusion criteria to extract high-quality issues. (i) The issue has been closed with a developer-committed patch to address the issue, as patches can serve as ground truth for understanding root causes of agent issues; (ii) The issue has clear descriptions without misleading information (e.g., exact patches or misleading patches in the problem description). This criteria has been widely used in constructing high-quality issue resolution benchmarks for traditional software systems DBLP:journals/corr/abs-2407-01489 ; sweverified ; swebench ; (iii) The issue should only report one problem instead of mixing multiple problems. In the end, we obtain 201 issues in total.
3.1.2 Manual Labeling
We randomly separate our collected 201 agent issues into (i) 171 issues (85%) for building the taxonomy and (ii) 30 issues (15%) for evaluating our constructed taxonomy.
Taxonomy construction. We manually catalog the 171 agent issues with grounded theory glaser2017discovery . In particular, three human annotators with extensive software development and machine learning experience apply open coding seaman1999qualitative to annotate each issue based on the issue description and the developer-committed patch. They break down each issue into segments and label them with descriptive codes. Then they organize the open codes into structured categories by merging and linking relevant ones. All the annotators further discuss and review the taxonomy until reaching a consensus.
Taxonomy evaluation. We further evaluate our taxonomy on the remaining 30 agent issues. Two annotators independently label each issue. Their annotation reaches a high agreement ratio (Cohen’s Kappa = 0.849); meanwhile there emerge no new categories in addition to our taxonomy during their annotation. The results suggest the generalizability and reliability of our taxonomy.
Table 1: Taxonomy of agent issues.
Category Sub-category Description Incompati- bility with LLM providers (7.46%) Incompatible dependencies (1.99%) Miss or misuse the libraries of LLM providers. Unsupported models (2.99%) Lack the support of recent LLMs. Incompatible parameters to LLM providers (2.49%) Use undefined parameters or miss parameters of LLM query interfaces. Tool-related issues (18.41%) Tool dependency issues (3.48%) Miss or misuse libraries for running agent-invoked tools. Tool configuration issues (2.49%) Misconfigure the settings of agent-invoked tools. Tool implementation errors (8.46%) Incorrect implementation of self-developed agent-invoked tools. Misuse tool interfaces (3.98%) Incorrect tool invocation due to missing/wrong parameters. Memory-related issues (14.43%) Memory initialization issues (2.49%) Incomplete or inconsistent memory states due to database initialization or workspace resetting issues. Memory content errors (10.95%) Incorrect message attributes, misleading content, or content loss caused by faulty storage logic. Memory dependency issues (1.00%) Incorrect internal module dependencies or external libraries required by memory operations. LLM operation issues (31.84%) Model access misconfiguration (7.46%) Model access errors caused by misconfiguration like incorrect model binding or authentication credentials (e.g., API keys). Token usage misconfiguration (3.48%) LLM token management issues such as incorrect limits or pricing. Incorrect model output handlers (8.46%) Incorrect parsing logic for model output or miss handlers for unexpected model behaviors like empty or exceptional responses. Model dependency issues (2.99%) Missing/incompatible libraries related to model operation such as tokenization or transformer dependency conflicts. Context length issues (4.98%) Truncated outputs caused by exceeding context limits or miscalculating context length. Prompt-related issues (4.48%) Suboptimal prompt content or prompt management issues (e.g., fail to set/update prompts). Workflow issues (6.97%) Abnormal agent workflows like hanging or repeated loops. Utility issues (20.90%) Utility implementation issues (9.45%) Implementation errors in LLM-unrelated components (e.g., UI/Docker/logging). Utility dependency issues (4.48%) Miss/incompatible libraries or circular internal dependencies required by general utilities (e.g., testing or file operations). Utility configuration issues (6.97%) External component misconfiguration (e.g., I/O paths, network settings).
3.2 Taxonomy
Table 1 presents our taxonomy of agent issues, mainly covering 6 categories. Appendix  F presents detailed examples for each sub-category. In addition to the “utility issues” category which may also occur in traditional software systems, the remaining five categories are uniquely tied to key agent system components (e.g., tools and memory), making them distinctive to agent systems.
Incompatibility with LLM providers. Most agent systems incorporate existing LLMs from LLM providers (e.g., OpenAI openai , DeepSeek deepseek , and Anthropic anthropic ), and improper usage of providers’ interfaces impairs agent functionality. Such issues often stem from missing dependencies or incorrect invocations of provider APIs. Moreover, due to the rapid evolution of LLMs, users frequently request new feature to support newly-released LLMs.
Tool-related issues. The versatility of agent systems partly stems from their proficiency in utilizing tools to interact with the environment. As a result, many agent-related issues arise during tool invocation, including missing tool-dependent libraries, misconfigurations, or incorrect use of tool interfaces. In addition to external tools, agents may also rely on internal tools (e.g., custom-developed functions), where implementation flaws can trigger unintended behaviors during tool execution.
Memory-related issues. The memory mechanism in agents tracks the trajectory of agent operation, and most memory-related issues arise from incorrect memory content. For example, agents may pollute memory with irrelevant information when they mistakenly extract unrelated attributes from the current context, or memory entries may be missing or incomplete due to failures in storing data.
Workflow issues. Due to the autonomy and flexibility of agent systems, unexpected behaviors can emerge along the agent workflow, such as repeated actions or hanging states. Although it is difficult to completely eliminate such issues, developers commonly mitigate them by incorporating status checkers to monitor and regulate the agent workflow.
LLM operation issues. A large portion (31.84%) of agent-related issues occur during LLM operation. For example, proper configuration of model access and token usage is critical, and misconfiguration in these areas can disrupt agent functionality. Additionally, many issues stem from incorrect handling of model outputs, including: (i) flawed parsing implementations, or (ii) missing handlers for unexpected model responses. Beyond the suboptimal prompt content (e.g., unclear model instructions), prompt management can also introduce risks: as agent systems often maintain a large and evolving pool of prompts, failures in prompt updates or configuration can result in models being queried with incorrect or outdated instructions.
Summary. Our taxonomy reveals that real-world agent systems exhibit a diverse range of issues, many of which possess unique characteristics not typically found in traditional software systems. In particular, developing and maintaining agent systems demands substantial engineering effort, as developers must manage correct dependencies, configurations, and implementations across multiple components (e.g., model providers, LLM operations, memory mechanisms, and tools). Therefore, we believe that automatically resolving issues in agent systems represents a challenging and increasingly vital research direction in the era of LLMs.
4 AgentIssue-Bench Benchmark
We then manually build AgentIssue-Bench, the first reproducible issue resolution benchmark of real-world agent issues. AgentIssue-Bench can be used to evaluate the efficacy of state-of-the-art SE agents in solving issues in agent systems.
4.1 Benchmark Construction
We construct AgentIssue-Bench out of the 201 GitHub agent issues we collected in Section 3. In particular, we try to reproduce each issue according to the following procedure.
Step 1: Failure reproduction. For each issue, we pull its corresponding buggy commit and set up the agent system. In particular, we manually write a test script (i.e., failure-triggering test) to reproduce the problematic behaviors according to the issue descriptions. In this step, we filter out the issues where we cannot observe the same buggy behavior as issue descriptions.
Step 2: Patch reproduction. We then pull the corresponding patched commit and execute the failure-triggering test on it. In this step, we only keep the issues where the patched version can pass the failure-triggering tests (i.e., problematic behaviors disappear on the patched version).
Step 3: Non-flakiness verification. Given the nondeterminism of LLMs, we repeat the previous two steps three times for each issue so as to eliminate the test flakiness. In this step, we filter out issues where there are inconsistent behaviors on executing one failure-triggering test.
Through such a multi-step filtering process, the original 201 agent issues are narrowed down to 50 reproducible issue resolution tasks, collectively forming AgentIssue-Bench. We find that reproducing issues in agent systems is significantly more challenging than in traditional software systems, as agent issues are associated with diverse internal and external components and resources. In particular, most agent issues fail to reproduce for the following reasons. (i) The nondeterminism of LLMs leads to unstable model outputs, which hinders the reproduction of agent issues such as workflow errors; (ii) External resources (e.g., agent-invoked tools, dependent libraries, or LLM providers) may have changed since the issue was reported, making it impossible to reproduce the same failure; (iii) Issue descriptions lack sufficient details or steps on how to reproduce the problematic behaviors; (iv) Agent systems cannot be correctly set up and exhibit unexpected failure behaviors that are different from the issue descriptions. Overall, the entire reproduction process takes huge manual effort (approximately 500 person-hours).
Figure 2: Distribution of AgentIssue-Bench
4.2 Benchmark Details
Benchmark statistics. Figure 2 shows the distribution of AgentIssue-Bench across different issue categories. Overall, we can observe that the 50 reproduced agent issues in AgentIssue-Bench cover all the main categories identified in our taxonomy of agent issues, indicating that AgentIssue-Bench is representative of real-world agent issue distribution. Moreover, issues in AgentIssue-Bench involve patches of different scales (Detailed statistics are in Table 5).
Each issue resolution instance in AgentIssue-Bench consists of the following components: (i) Issue description: a user-reported textual description of the problem; (ii) Buggy version of the agent system: the buggy commit of the agent code repository in which the issue occurs; (iii) Developer-committed patch: the code changes between the buggy and correct versions, serving as the ground truth for issue resolution; (iv) Failure-triggering tests: test scripts that reproduce the issue on the buggy version but pass on the patched version; (v) Docker environment: a container with all necessary dependencies and configurations to execute the agent system.
Figure 3: A task example in AgentIssue-Bench.
Task formulation. The agent issue resolution task can be formulated as follows: (i) Input: the issue description and the buggy codebase of the agent system; (ii) Output: a patch (i.e., a code edit to the buggy codebase) that aims to resolve the issue. Figure 3 shows the task example in AgentIssue-Bench.
Evaluation metrics. To evaluate how a technique tackles the agent issue resolution task, we adopt the following metrics to evaluate the patches output by the technique (i.e., SE agents in our experiments). (i) Localization accuracy: if the generated patch modifies the same location as the developer-committed patch, we consider it to have accurately localized the issue. We then compute the percentage of issues for which the generated patches can achieve accurate localization. (ii) Plausible resolution rate: if the generated patch makes the failure-triggering tests pass after being applied, we consider it to plausibly resolve the issue (i.e., denoted as a plausible patch). We then compute the percentage of issues for which the generated patches are plausible patches. (iii) Correct resolution rate: if the generated plausible patch is further semantically-equivalent to the developer-committed patch, we consider it to correctly resolve the issue (i.e., denoted as a correct patch). In particular, given the insufficiency of tests in practice, it is common 10.1145/3663529.3663776 ; 9034821 ; 10.1145/2884781.2884872 that plausible patches are not necessarily correct patches but are just overfitting to the failure-triggering tests. Therefore, only reporting the plausible resolution rate can overestimate the effectiveness of issue resolution techniques. Following the common practice in the program repair area chatrepair ; alpharepair ; xiaase23 ; knod , we further involve human annotators to manually check whether the plausible patches are semantically equivalent to developer-committed patches. We then compute the percentage of issues for which the generated patches are correct patches.
5 Experiments
In this section, we investigate how state-of-the-art SE agents can automatically resolve real-world issues in agent systems by evaluating their efficacy on AgentIssue-Bench.
5.1 Experimental Setup
Studied SE agents. We include three state-of-the-art SE agents, including SWE-agent DBLP:journals/corr/abs-2405-15793 , AutoCodeRover DBLP:conf/issta/0002RFR24 , and Agentless DBLP:journals/corr/abs-2407-01489 . These agents are selected given that they are fully open-sourced and achieve superior effectiveness in resolving issues for traditional software systems sweboard . We directly adopt their released implementation with the original hyperparameter settings.
Backbone LLMs. Based on the recent SWE leaderboard sweboard , state-of-the-art SE agents achieve higher fixing rate on general software issues when equipped with backbone LLMs GPT-4o GPT4o and Claude-3.5 Sonnet claude . Therefore, in our experiments, we mainly study how effective SE agents are in resolving agent issues with these two backbone LLMs.
Evaluation pipelines. We apply studied SE agents on AgentIssue-Bench and collect their generated patches for each issue resolution task. We then calculate the metrics of fault localization accuracy, plausible and correct resolution rates for each studied SE agent. To eliminate the randomness from LLMs, we repeat all experiments three times and present the average results.
Table 2: Overall results of SE agents on AgentIssue-Bench
SE Agent LLM Plausibly resolved% Correctly resolved% Localization % Avg. File-level Function-level \$Cost Agentless GPT-4o 18.67 6.00 22.97 14.40 0.65 Claude-3.5-S 12.67 8.67 21.86 11.54 0.33 AutoCodeRover GPT-4o 12.67 4.67 17.05 11.30 0.23 Claude-3.5-S 17.33 12.67 25.61 18.67 0.05 SWE-agent GPT-4o 5.33 3.33 12.65 11.77 1.15 Claude-3.5-S 6.67 6.67 15.58 11.26 0.57
5.2 Quantitative Results
Overall resolution effectivenes. Table 2 shows the results of the studied SE agents on AgentIssue-Bench. In general, state-of-the-art SE agents can only correctly resolve a small number (i.e., 3.33% - 12.67%) of agent issues. In addition, in most cases, SE agents even fail to correctly identify the location (i.e., files or functions) for resolving the issue, e.g., file-level/function-level localization accuracy is less than 26%/19%. Such observations reveal the limited capabilities of state-of-the-art SE agents in understanding and resolving the issues in agent systems.
In addition, Figure 4 compares the correct resolution rate of SE agents on agent issues (on our benchmark AgentIssue-Bench) versus on traditional software issues (results from SWE-bench Lite sweboard ). As there is no previous data of AutoCodeRover with Claude-3.5-S on SWE-bench, we leave it as blank. Overall, SE agents demonstrate significantly lower resolution rates on agent issues compared to traditional software issues. These findings highlight the unique challenges posed by agent systems and underscore the need for developing SE agents specifically tailored to maintain agent systems, which is an emerging and distinctive software paradigm.
Comparison among SE agents and backbone LLMs. As shown in Table 2, SE agents with Claude-3.5-S achieve higher resolution capabilities than with GPT-4o in terms of plausible resolution, correct resolution, and localization accuracy. In particular, AutoCodeRover with Claude-3.5-S achieves the highest resolved rate (i.e., 12.67%) and the highest localization accuracy (i.e., 25.61% at file level). Overall, we observe a larger potential of Claude-3.5-S in understanding agent issues than GPT-4o.
Figure 5 shows the unique and overlapped agent issues that are correctly resolved by each SE agent. We could observe that each SE agent can uniquely fix 2 - 4 bugs that cannot be resolved by any other SE agents. In addition, there is no agent issue that can be fixed by all SE agents. In other words, existing SE agents exhibit complementary capabilities to resolve agent issues.
Figure 4: Resolution rate of agent issues v.s. traditional software issues.
Figure 5: Venn diagrams of resolved issues.
Costs. As shown in Table 2, the average costs of applying SE agents to agent issue are controllable, ranging from \$0.05 to \$1.15. The cost range is similar as applying these SE agents to resolve traditional software issues (e.g., \$0.45 - \$2.53 DBLP:journals/corr/abs-2407-01489 ).
Table 3: Breakdown of resolved agent issues (unresolved categories are not presented).
Category Resolved% Sub-category Resolved% Tool-related issues 3/9 (33.33%) Tool dependency issues 2/3 (66.67%) Tool implementation issues 1/3 (33.33%) Memory-related issues 1/8 (12.50%) Memory content errors 1/5 (20.00%) LLM operation issues 2/11 (18.18%) Model access misconfiguration 1/2 (50.00%) Prompt-related issues 1/2 (50.00%) Workflow issues 1/6 (16.67%) - - Utility issues 8/14 (57.14%) Utility implementation issues 2/6 (33.33%) Utility dependency issues 2/2 (100.00%) Utility configuration issues 4/6 (66.67%)
5.3 Qualitative Results
In this section, we further break down the issues that SE agents can and cannot resolve, aiming to better understand their strengths and limitations in resolving agent issues. Table 3 presents the issue categories that can be resolved by at least one studied SE agent.
Resolved agent issues. Overall, the majority of agent issues resolved by SE agents are still related to utility (e.g., log/file operation/UI), which actually share high commonality with traditional software systems. As a result, SE agents are inherently able to resolve issues of this category in agent systems. Moreover, besides common utility issues, some of the dependency issues on agent-specific components (e.g., tool) can also be resolved by SE agents. The reason why SE agents can handle such agent issues might be that the dependency issues often contain explicit error messages (e.g., missing libraries or incompatible variables/interfaces). As a result, even if the dependencies are unique to agent components (e.g., tool), they can still be similar to dependency issues in other general software components, which are straightforward and informative to resolve.
Unresolved agent issues. Overall, the majority of agent-specific issues cannot be resolved by any SE agent. For example, SE agents resolve a very few (or even none) issues on LLM provider incompatibility, memory, or LLM operation. The reason might be that the exchanges with LLM providers are unique features in agent systems and agent systems are emerging in the recent period, which thus are less covered in the LLM training data. In addition, the autonomous and flexible nature of agent systems stemming from LLMs makes it challenging to identify the root causes of LLM operation issues. Figure 6 and Figure 7 in Appendix E show two unresolved issues for which all SE agents cannot even correctly localize the buggy files.
In summary, our analysis further confirms the limitations of existing SE agents in resolving the agent issues which are particularly related to agent-specific features, highlighting the necessity of building more advanced SE agents for maintaining agent systems.
6 Limitation and Future work
While AgentIssue-Bench is representative of real-world agent issues by covering a wide range of different categories, the generality of our findings can still be restricted due to the current size of AgentIssue-Bench. In particular, we find reproducing issues in agent systems is significantly more challenging than in traditional software systems. Due to the nondeterminism of LLMs and changeable external resources (e.g., tools and LLM providers) interacted with agent systems, only a small number of agent issues (50 out of 201 issues) can be successfully reproduced. Moreover, huge manual effort (approximately 500 person-hours) are dedicated to preparing the Docker environment, configuring agent systems, and writing failure-triggering tests. In the future, we plan to continuously maintain and extend the benchmark to support future research on agent system maintenance.
7 Conclusion
In this work, we analyze 201 GitHub issues from 16 real-world agent systems and construct the first taxonomy of agent issues. We further build AgentIssue-Bench, the first reproducible benchmark of 50 high-quality agent issue resolution tasks. Experiments on state-of-the-art SE agents demonstrate their limited effectiveness in addressing agent issues (with resolution rates ranging from 3.33% to 12.67%), highlighting the unique challenges in maintaining agent systems and the pressing need for more advanced SE agents tailored to this emerging software paradigm.
Appendix A Analyzed agent systems
Table 4 presents the full list of our analyzed agent systems and their statistics.
Table 4: Statistics of analyzed agent systems
Agent #stars #Loc Creation time License agent-squad [4] 5.4k 109,019 07/23/2024 Apache-2.0 License AGiXT [5] 3k 111,946 04/04/2023 MIT License autogen [8] 44.2k 197,969 08/18/2023 CC-BY-4.0, MIT License camel [10] 12.4k 206,152 03/17/2023 Apache-2.0 License babyagi [9] 21.4k 8,800 04/03/2023 MIT License CrewAI [13] 31.3k 171,395 10/27/2023 MIT License Lagent [21] 2.1k 13,075 08/20/2023 Apache-2.0 License MetaGPT [22] 55.4k 90,709 06/30/2023 MIT License RagaAI-Catalyst [26] 16.2k 47,252 10/01/2024 Apache-2.0 License ChatDev [11] 26.8k 40,478 11/04/2023 Apache-2.0 License gpt-engineer [18] 54.1k 17,460 04/29/2023 MIT License Pythagora [25] 1.8k 5,859 01/21/2023 Apache-2.0 License SWE-agent [28] 15.7k 63,388 04/02/2024 MIT License evo.ninja [16] 1.1k 31,862 08/18/2023 MIT License Superagent [27] 5.8k 58,602 05/10/2023 MIT License gpt-researcher [19] 21.3k 168,849 05/12/2023 Apache-2.0 License
Table 5: Mean and maximum values for various patch attributes in studied agents
Attribute Mean Max # Lines edited 66.05 355 # Files edited 3.58 34 # Functions edited 6.79 54
Appendix B Patch Scales of AgentIssue-Bench
Table 5 presents the statistics of the patch scales in AgentIssue-Bench.
Appendix C Human Participation
All human-involved tasks in our experiments (including taxonomy construction, taxonomy evaluation, and issue reproduction) were approved by the Institutional Review Board (IRB) at our institution. Additionally, all participants were compensated at a rate of \$15 per hour.
In taxonomy evaluation, each human annotator is provided with the following instructions: “Given the taxonomy of agent issues (along with the definition of each agent issue category), please label each agent issue with any category in the taxonomy. If there are no applicable categories in the given taxonomy, please return the label as non-applicable.”
Appendix D Experiment Statistical Significance
Table  6 presents the mean results of SE agents on AgentIssue-Bench, along with their corresponding two-sigma ($\pm{}2\sigma$) errors. To calculate the two-sigma errors, we conducted the experiment on AgentIssue-Bench three times, computed the standard deviation for the results and multiplied it by 2, as shown in Equation 1:
$2\sigma=2\times\sqrt{\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}}$ (1)
where $N$ is the number of experimental runs, $x_{i}$ is the result of the $i$-th run, and $\bar{x}$ is the mean result.
Table 6: Mean results with 2-sigma($\pm{}2\sigma$) errors of SE agents on AgentIssue-Bench
SE Agent LLM Plausibly resolved% Correctly resolved% Localization % File-level Function-level Agentless GPT-4o 18.67 ($\pm{}2.31$) 6.00 ($\pm{}4.00$) 22.97 ($\pm{}4.74$) 14.40 ($\pm{}3.23$) Claude-3.5-S 12.67 ($\pm{}2.31$) 8.67 ($\pm{}2.31$) 21.86 ($\pm{}3.81$) 11.54 ($\pm{}1.36$) AutoCodeRover GPT-4o 12.67 ($\pm{}4.62$) 4.67 ($\pm{}4.62$) 17.05 ($\pm{}1.79$) 11.30 ($\pm{}1.03$) Claude-3.5-S 17.33 ($\pm{}4.62$) 12.67 ($\pm{}2.31$) 25.61 ($\pm{}3.18$) 18.67 ($\pm{}6.67$) SWE-agent GPT-4o 5.33 ($\pm{}4.62$) 3.33 ($\pm{}4.62$) 12.65 ($\pm{}5.06$) 11.77 ($\pm{}6.94$) Claude-3.5-S 6.67 ($\pm{}2.31$) 6.67 ($\pm{}2.31$) 15.58 ($\pm{}7.00$) 11.26 ($\pm{}1.59$)
Appendix E Examples of Unresolved Issues
In this section, we provide two issue examples that all SE agents fail to localize the buggy files and generate correct patches.
For the example depicted in Figure 6, the agent lacks up-to-date knowledge regarding which LLMs currently support the “stop” parameter. As a result, the agent incorrectly passed the “stop” parameter to LLMs that do not support it (i.e., o1-preview and o1-mini), ultimately aggravating the issue.
For the example depicted in Figure 7, the agent fails to identify the root cause of the KeyError, i.e., a conflict arising from generating multiple diffs for a single file. This issue is specific to the agent system, as it involves the handling of model outputs. However, instead of performing a deeper analysis, the agent merely prints a error message, resulting in an unsuccessful patch.
Figure 6: This unresolved issue arises because not all LLMs support the ‘stop’ parameter, requiring users to control its use (e.g., via use_stop_words in the Golden Patch). The agent-generated patch aggravated the issue by passing ‘stop’ to unsupported models (i.e., o1-preview and o1-mini).
Figure 7: This unresolved issue stems from the LLM generating multiple diffs for a single file, resulting in conflicts. The Golden Patch resolves this by retaining only the first diff. In contrast, the agent-generated patch fails to investigate the root cause and simply logs an error message.
Appendix F Examples of Issues in Different Categories
In this section, we provide detailed issue examples of each sub-category in Table 1. For each issue, we provide the repository name, the user-provided issue description and the summarization of the developer-committed patch (along with the link to the original issue and Pull Request (PR) pages).
F.1 Incompatibility with LLM providers
  1. 1.
    Incompatible dependencies
  2. 2.
    Unsupported models
  3. 3.
    Incompatible parameters to LLM providers
    • Repository: CrewAI
    • Issue Description: I defined the model to o1-preview or o1-mini and temperature to 1, and I get the following error. Unsupported parameter: “stop” is not supported with this model. Apparently the stop parameter is used, and is not supported. I didn’t find a way for the crew to no use this parameter.
    • Fix Strategy: Added the option use_stop_words to allow users to configure whether to use the stop parameter.
F.2 Tool-related issues
  1. 1.
    Tool dependency issues
  2. 2.
    Tool configuration issues
  3. 3.
    Tool implementation issues
  4. 4.
    Misuse tool interfaces
    • Repository: camel
    • Issue Description: When “KAUST” is the entity word to be searched, the returned result by wikipedia API (wikipedia.summary) is the summary about KAIST.
    • Fix Strategy: Set the auto_sugget parameter to False when invoking wikipedia.summary() so that it does not change the search word (such as KAUST -> KAIST)
F.3 Memory-related issues
  1. 1.
    Memory initialization issues
  2. 2.
    Memory content issues
    • Repository: camel
    • Issue Description: Current (memory storage) logic checks the content in the chunk, if the content is None then the message would be appended, but for some API like SambaNova, there may include many None content in chunks in the middle of response. We need to change the logic, checking choice.finish_reason then append the message would be better.
    • Fix Strategy: Update the logic for storing messages. Determine whether a chunked message has been fully stored by checking if the finish_reason attribute is not None.
  3. 3.
    Memory dependency issues
    • Repository: autogen
    • Issue Description: Running autogenstudio ui –port 8081 fails with ImportError: cannot import name ‘InnerMessage’ from ‘autogen_agentchat.messages’
    • Fix Strategy: Since ‘InnerMessage’ has been renamed to ‘AgentMessage’, all references to ‘InnerMessage’ are renamed to ‘AgentMessage’.
F.4 LLM operation issues
  1. 1.
    Model access misconfiguration
    • Repository: camel
    • Issue Description: The decorator api_keys_required() currently only supports setting the value of “DUMMY_TOKEN” in the environment variable, but does not support directly calling DummyClass(api_key="xxxx").
    • Fix Strategy: Refactor the api_keys_required() decorator and make it compatible with the method of directly setting API key.
  2. 2.
    Token usage misconfiguration
    • Repository: camel
    • Issue Description: For LLM API served by OpenAI-compatible providers, if the max_tokens is not provided then it would use NOT_GIVEN from openai, which will lead to TypeError: ‘<=’ not supported between instances of ‘int’ and ‘NotGiven’
    • Fix Strategy: Unify the default token_limit property in the base model class to make sure it is provided for different models.
  3. 3.
    Incorrect model output handlers
    • Repository: MetaGPT
    • Issue Description: When running python3 debate.py “Talk about Artificial General Intelligence”, an error occurs: ValueError: The response.text quick accessor only works for simple (single-Part) text responses.
    • Fix Strategy: The root cause is that the Gemini model flags the request as potentially involving sensitive or harmful content and blocks it. To address this scenario, the BlockedPromptException has been added to catch exceptions triggered by blocked prompts.
  4. 4.
    Model dependency issues
  5. 5.
    Context length issues
  6. 6.
    Prompt-related issues
F.5 Workflow issues
  1. 1.
    Workflow issues
F.6 Utility issues
  1. 1.
    Utility implementation issues
    • Repository: SWE-agent
    • Issue Description: In the React frontend framework, the default value of a textbox is supposed to update based on another selected dropdown item, but this dynamic binding is not functioning as intended.
    • Fix Strategy: Use useEffect to monitor the change of the dropdown item, and when it changes, reset the textbox to display the default value.
  2. 2.
    Utility dependency issues
  3. 3.
    Utility configuration issues
    • Repository: CrewAI
    • Issue Description: When loading a variable from the yaml environment, the crewai library tries to load it in gbk encoding as soon as Chinese is present in it, but my yaml file is utf-8. You want to add a measure to detect file encoding when loading yaml files.
    • Fix Strategy: Explicitly specify UTF-8 encoding when reading the YAML file.